how much pure alcohol must be added to 4 liters of 10% alcohol solution to make a solution that is 20% alcohol?
Answer by TimothyLamb(4379)

---
10*4 + 100v = 20(v + 4)
40 + 100v = 20v + 80
80v = 40
v = 0.5
---
answer:
add 0.5 liter of pure alcohol, and you'll have 4.5 liters of 20% alcohol
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php